# Quantum variational calculus on a lattice

### v2

@@ 1. Introduction
(1) $E[\phi]=-{1\over 2}(\partial_{+}\phi)(\partial_{-}\phi)+{m^{2}\over 2}\phi^{2}$
@@ 2.1. Jet bundle and variational bicomplex
As a check, we see that $\mathrm{d}=\mathrm{d}_{H}+\mathrm{d}_{V}$ holds on $C^{\infty}(J^{\infty})$. This induces a factorisation of the deRham complex into a double complex
@@
Note that it does not matter too much what $\mathrm{d}_{F}\partial_{I}$ is as it cancels in the calculation, but we can let $\phi=\int\delta_{y}\phi^{y}\mathrm{d}y$ be an expansion in a delta-function $\delta_{y}(x)$ basis and coordinatize $F$ by the pointwise values $\{\phi^{y}\ |\ y\in\mathbb{R}$}. Then if $\Psi\in C^{\infty}(F)$, $\mathrm{d}_{F}\Psi=\int\mathrm{d}y{\partial\Psi\over\ \partial\phi^{y}}\mathrm{d}\phi^{y}$ and hence in particular,
$e_{\infty}^{*}(\mathrm{d}u_{I})=(\partial_{iI}(\cdot))(\cdot)\mathrm{d}x^{i}+\int\mathrm{d}y\left({\partial\over\partial\phi^{y}}\partial_{I}(\cdot)\right)(\cdot)\mathrm{d}\phi^{y}=(\partial_{iI}(\cdot))(\cdot)\mathrm{d}x^{i}+\int\mathrm{d}y(\partial_{I}\delta_{y})\mathrm{d}\phi^{y}$
where we assume the $\delta_{y}$ functions have been smoothed so that we can differentiate them as functions of $x$. The second term is constant on $F$ and we have also suppressed that it is a function on $M$. We do not have to chose a $\delta$-function basis and more reasonable here would be a plane-wave basis i.e. to coordinatize $\phi\in F$ by its Fourier coefficients. These matters can be made more precise by usual methods in mathematical physics but this is not needed for our purposes.
@@ 2.2. Euler-Lagrange equations
Computing the RHS gives $\mathrm{d}_{V}L\wedge\mathrm{Vol}=EL-\mathrm{d}_{H}\Theta$ for certain $EL\in\Omega^{{\mathrm{top}},1}(J^{\infty})$ the EL form and $\Theta\in\Omega^{{\mathrm{top}}-1,1}(J^{\infty})$ the boundary term. Explicitly, we have
@@
This result can be generalised to Lagrangians including higher derivative terms $u_{I}$ [1], but 1st order Lagrangians that are functions of one derivative are sufficient for us.
@@
Consider free scalar field theory on base $M=\mathbb{R}^{n}$ with the Euclidean metric or $M=\mathbb{R}^{1,n-1}$ with the Minkowski metric, both denoted by $g$, then the action is
@@ 2.3. Symmetries and Noether’s Theorem
For charges that are conserved in a time direction, consider $M=\mathbb{R}\times\Sigma$ for some codimension 1 submanifold $\Sigma$ without boundary, with $\mathbb{R}$ denoting the time direction. In this setting, the coordinate on $\mathbb{R}$ will be denoted with the index 0 and the coordinates on $\Sigma$ with the index $i$, so we write $j_{X}=j^{0}_{X}\mathrm{Vol}_{0}+\sum^{\dim\Sigma}_{i=1}j^{i}_{X}\mathrm{Vol}_{i}$. We continue to focus on the Euclidean and Minkowski signatures and identify $\mathrm{Vol}_{\Sigma}$ with the volume form on $\Sigma$.
@@
Consider $E=M\times\mathbb{R}\to M$, with $M=\mathbb{R}^{n}$ with the Euclidean or Minkowski metric and a first-order Lagrangian $L=L(u,u_{i})$ which is translation invariant. This symmetry can be encoded with $\iota_{X_{\infty}}(\mathrm{d}x^{i})=\epsilon^{i}$, $\iota_{X_{\infty}}(\mathrm{d}u_{I})=0$ so that $\iota_{X_{V}}(\mathrm{d}_{V}u_{I})=-\epsilon^{i}u_{iI}$. We have
@@
(4) $P_{i}[\phi]=\int_{\Sigma}T_{0i}[\phi]\,\mathrm{d}^{n-1}x=\int_{\Sigma}\partial_{0}\phi\partial_{i}\phi\,\mathrm{d}^{n-1}x.$
In particular the energy for the Euclidean and Minkowski metrics with $g_{00}=1$, $g_{ii}=\pm 1$ respectively is
@@
Focusing now on the (1+1)-case for simplicity and choosing a plane wave solution $\phi(t,x)=A\cos(\omega t+g_{11}\kappa x)$ solving the Klein-Gordon equation $(\Delta+m^{2})\phi=0$ with the dispersion relations $m^{2}=\omega^{2}+g_{11}\kappa^{2}$, one recovers that the expected energy and momentum densities (the integrands of $E[\phi]$ and $P_{x}[\phi]$) are given by
@@
and choosing the normalization $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ of the field $\phi$ we then recover the familiar $\langle\mathcal{E}_{dens}[\phi]\rangle^{2}+g_{11}\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}=m^{2}$, indicating that the field $\phi$ has the density of ‘one particle of mass $m$ per unit volume’, with $\langle\mathcal{E}_{dens}[\phi]\rangle=\omega$ and $\langle\mathcal{P}_{x,dens}[\phi]\rangle=\kappa$. Note that the usual normalization $\sqrt{1/(2\langle\mathcal{E}_{dens}[\phi]\rangle)}$ found when quantizing the field via the CCR is recovered if we write the field $\phi(t,x)=(A/2)(e^{{\mathfrak{i}}(\omega t+g_{11}\kappa x)}+e^{-{\mathfrak{i}}(\omega t+g_{11}\kappa x)})$ in terms of exponentials as it is done when quantizing it.
@@
Note that $\mathcal{E}_{dens}[\phi,\phi^{*}]$, $\mathcal{P}_{x,dens}[\phi,\phi^{*}]$ are double the values of the real scalar field theory in Example 2.6 due to the existence of two different modes $\phi$, $\phi^{*}$ in this theory. In this case we find the relations
@@ 3. Variational calculus on the integer lattice $\mathbb{Z}$
We now follow the exact same ideology but replacing the base manifold $M$ with $\mathbb{Z}$ as a discrete noncommutative geometry. This is an application of a general picture whereby $\Omega^{1}$ on a discrete set is precisely a graph, and here we consider $\mathbb{Z}$ as a graph, namely a one-dimensional lattice. Using that $\mathbb{Z}$ is a group, it is convenient to write the calculus in terms of a basis of invariant 1-forms $e^{a}=e^{\pm}$ over the algebra of real valued functions $C(\mathbb{Z})$. The exterior derivative is then
@@ 3.1. Construction of the double complex $\Omega(J^{\infty})$
with coordinates $(i,u_{I})$. The jet prolongation map is
@@
$e_{\infty}^{*}(\Phi)(i,\phi)=\Phi(i,(\partial_{I}\phi)(i)),\quad e_{\infty}^{*}(e^{\pm})=\sum_{i}e_{\infty}^{*}(\delta_{i}\mathrm{d}\delta_{i\pm 1})=e^{\pm}$
@@
and we recognising the last term, we obtain
@@
where the second relation is redundant i.e. can be proved by application of the other relations. We used that the $e^{a}$ anticommute to obtain it in this form. Next, using (8) and assuming $\mathrm{d}_{H}$ is a derivation as it should be as inherited from $\Omega(\mathbb{Z}\times F)$ and that $\mathrm{d}_{H}e^{a}=\mathrm{d}e^{a}=0$, we have
@@
$\mathrm{d}_{V}u_{I},\Phi]=0,\quad\{e^{a},\mathrm{d}_{V}u_{I}\}+\mathrm{d}_{V}u_{aI}\wedge e^{a}=0,\quad\{\mathrm{d}_{V}u_{I},\mathrm{d}_{V}u_{J}\}=0$
@@
$[\mathrm{d}_{V}u_{I},\phi]$ $=[\mathrm{d}u_{I},\phi]-\sum_{a}u_{aI}[e^{a},\phi]=[\mathrm{d}u_{I},\phi]-\sum_{a}u_{aI}(\partial_{a}\phi)e^{a}=0$
$[\mathrm{d}_{V}u_{I},u_{J}]$ $=[\mathrm{d}u_{I},u_{J}]-\sum_{a}u_{aI}[e^{a},u_{J}]=0$
@@ 3.2. Euler-Lagrange equations
$R_{a}\colon\Omega^{k}(J^{\infty})\to\Omega^{k}(J^{\infty})\qquad D_{a}\colon\Omega^{k}(J^{\infty})\to\Omega^{k}(J^{\infty})$
@@
(13) $R_{a}(\omega\wedge\eta)=R_{a}(\omega)\wedge R_{a}(\eta)\qquad D_{a}(\omega\wedge\eta)=D_{a}(\omega)\wedge R_{a}\eta+\omega\wedge D_{a}(\eta).$
@@
$EL=\left(\frac{\partial L}{\partial u}+D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\right)\mathrm{d}_{V}u\wedge\mathrm{Vol},\qquad\Theta=R_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\wedge\mathrm{Vol}_{a}.$
@@
$\mathrm{d}_{V}(L\mathrm{Vol})=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u_{a}\right)\wedge\mathrm{Vol}=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\frac{\partial L}{\partial u_{a}}D_{a}\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}$
@@
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+D_{a}\left(\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u\right)-D_{a}\left(\frac{\partial L}{\partial u_{a}}\right)R_{a}\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}$
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+D_{a}\left(\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u\right)+R_{a}\left(D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)\right)\wedge\mathrm{Vol}$
@@
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+D_{a}\left(\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u\right)+D_{a}\left(D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)+D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}.$
@@
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+D_{a}\left(R_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)+D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}.$
@@
$\mathrm{d}_{V}(L\mathrm{Vol})=\left(\frac{\partial L}{\partial u}+D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\right)\mathrm{d}_{V}u\wedge\mathrm{Vol}-\mathrm{d}_{H}\left(R_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\wedge\mathrm{Vol}_{a}\right).$
@@
As the bundle $\mathbb{Z}\times\mathbb{R}\to\mathbb{Z}$ models a point particle moving in $\mathbb{R}$ evolving in discrete time, we want to reproduce Example 2.1 in this setting, starting with a free particle $q=\phi:\mathbb{Z}\to\mathbb{R}$ with $V(q)=0$. Let $(e^{a},e^{b})=g^{ab}=\delta^{a,b^{-1}}$ represent the inverse Euclidean metric on $\mathbb{Z}$, and define the integral on $\mathbb{Z}$ as $\int_{\mathbb{Z}}f\mathrm{Vol}=\sum_{i}f_{i}$. Then in analogy to Example 2.1 we set the action to be
@@
From the general discussion we take the key requirement of an interior product along a suitable vector field enconding the translation symmetry and we specify this directly as a map $\iota_{\epsilon}\colon\Omega^{\bullet}(J^{\infty})\to\Omega^{\bullet-1}(J^{\infty})$ by defining it on $\Lambda_{J^{\infty}}$ and extending as a left module map to $\Omega(J^{\infty})=C(J^{\infty})\Lambda_{J^{\infty}}$. Following the treatment for $M$ in Example 2.6, we let
@@
$=\sum_{b}\epsilon^{b}u_{abI}e^{a}=\sum_{b}\epsilon^{b}u_{baI}e^{a}=-\iota_{V}(\mathrm{d}_{V}u_{aI}e^{a})$
@@
(14) $i_{H}({\rm Vol})=i_{H}(e^{+}\wedge e^{-})=\epsilon^{+}e^{-}-e^{+}\epsilon^{-}=\sum_{a}\epsilon^{a}{\rm Vol}_{a}$
@@
where (,)(\ ,\ ) is the metric inner product and in our case $\nabla$ obeys $\nabla e^{a}=0$ and the Leibniz rule and extends to tensor products in an obvious way (it is a bimodule connection in the sense used in [2] with the generalised braiding given by flip on the basis). Hence divergence-free just amounts to $\sum_{a}\partial_{a^{-1}}(T_{ab})=0$ for all $b$.
@@
$\sum_{a}D_{a^{-1}}T_{ab^{-1}}=-{m\over 4}\sum_{a}\left(2u_{aa^{-1}}+D_{b^{-1}}\left(u_{aa^{-1}}\right)\right)u_{b^{-1}}$
We therefore see that the last term in the changed expression for $\sum_{a}D_{a^{-1}}T_{ab}$ can be writen in terms of $\sum_{a}D_{a^{-1}}T_{ab^{-1}}$, leading to
@@
where we have used $D_{a}=-D_{a^{-1}}R_{a}$, and directly see see that the term in the brackets is preserved along discrete time $\mathbb{Z}$. Expanding it out gives
@@
$E[q]=-{m\over 2}(\partial_{+}q)(\partial_{-}q)$
@@
The current from Proposition 3.4 is conserved when a quadractic potential $V(q)=\mu q^{2}$ is included in the Lagrangian $L(u,u_{a})={m\over 4}\sum_{a}u^{2}_{a}-V(u)$.
@@
$\mathrm{d}_{H}j=-{m\over 4}\sum_{a,b}\epsilon^{b}\left(2u_{aa^{-1}}u_{b}+D_{b}(u_{aa^{-1}})u_{b}\right)+\sum_{b}\epsilon^{b}(D_{b}V(u))\mathrm{Vol}.$
@@
Figure 1. Discrete time harmonic oscillator for $m=1$ and energy $E=1$ with increasing $\omega$. For the sine modes with $\omega<2$, the dashed line is the amplitude $q_{2}$ such that $E={1\over 8}m\omega^{2}(4-\omega^{2})q_{2}^{2}$, with minimum amplitude for $\omega=\sqrt{2}$. The period $2\pi/\omega^{\prime}$ is significantly smaller than $2\pi/\omega$ as $\omega$ increases. As $\omega\to 2$ from below, the amplitude blows up and the period tends towards the minimum alternating case. We also see a beat frequency emerging. The $\omega>2$ case has alternating exponential growth. Plots have been smoothly interpolated for visualisation.

### v3

@@ 1. Introduction
$E[\phi]=-{1\over 2}(\partial_{+}\phi)(\partial_{-}\phi)+{m^{2}\over 2}\phi^{2}$
@@ 2.1. Jet bundle and variational bicomplex
Here, $\mathrm{d}_{V}u_{I}=\mathrm{d}u-u_{iI}\mathrm{d}x^{i}$ are called contact 1-forms. As a check, we see that $\mathrm{d}=\mathrm{d}_{H}+\mathrm{d}_{V}$ holds on $C^{\infty}(J^{\infty})$. This induces a factorisation of the deRham complex into a double complex
@@
Note that it does not matter too much what $\mathrm{d}_{F}\partial_{I}$ is as it cancels in the calculation, but we can let $\phi=\int\delta_{y}\phi^{y}\mathrm{d}y$ be an expansion in a delta-function $\delta_{y}(x)$ basis and coordinatise $F$ by the pointwise values $\{\phi^{y}\ |\ y\in\mathbb{R}$}. Then if $\Psi\in C^{\infty}(F)$, $\mathrm{d}_{F}\Psi=\int\mathrm{d}y{\partial\Psi\over\ \partial\phi^{y}}\mathrm{d}\phi^{y}$ and hence in particular,
$e_{\infty}^{*}(\mathrm{d}u_{I})=(\partial_{iI}(\cdot))(\cdot)\mathrm{d}x^{i}+\int\mathrm{d}y\left({\partial\over\partial\phi^{y}}\partial_{I}(\cdot)\right)(\cdot)\mathrm{d}\phi^{y}=(\partial_{iI}(\cdot))(\cdot)\mathrm{d}x^{i}+\int\mathrm{d}y(\partial_{I}\delta_{y})(\cdot)\mathrm{d}\phi^{y}$
where we assume the $\delta_{y}$ functions have been smoothed so that we can differentiate them as functions of $x$. The second term is constant on $F$ and we have also suppressed that it is a function on $M$. We do not have to chose a $\delta$-function basis and more reasonable here would be a plane-wave basis i.e. to coordinatise $\phi\in F$ by its Fourier coefficients. These matters can be made more precise by usual methods in mathematical physics but this is not needed for our purposes.
@@ 2.2. Euler-Lagrange equations
Computing the RHS gives $\mathrm{d}_{V}L\wedge\mathrm{Vol}=EL-\mathrm{d}_{H}\Theta$ for certain $EL\in\Omega^{{\mathrm{top}},1}(J^{\infty})$ the EL form and $\Theta\in\Omega^{{\mathrm{top}}-1,1}(J^{\infty})$ the boundary term [30]. Explicitly, we have
@@
This result can be generalised to Lagrangians including higher derivative terms $u_{I}$ [1], but 1st order Lagrangians that are functions of one derivative are sufficient for us. We say that a field $\phi\in F$ satisfies the equations of motion, or is ‘on-shell’, if
(2) $e^{*}_{\infty}(EL)(\cdot,\phi)=0,$
or equivalently
$e^{*}_{\infty}\left(\frac{\partial L}{\partial u}-D_{i}\left(\frac{\partial L}{\partial u_{i}}\right)\right)(\cdot,\phi)=0.$
@@
Consider free scalar field theory on the base $M=\mathbb{R}^{n}$ with the Euclidean metric or $M=\mathbb{R}^{1,n-1}$ with the Minkowski metric, both denoted by $g$, then the action is
@@ 2.3. Symmetries and Noether’s Theorem
Here the current $j_{X}$ is conserved ‘on-shell’ in the sense that it is $\mathrm{d}_{H}$-closed up to the $EL$ form, i.e. $e^{*}_{\infty}(\mathrm{d}_{H}j_{X})(\cdot,\phi)=0$ for fields $\phi$ which satisfy equation (2). To consider charges that are conserved in a time direction, consider $M=\mathbb{R}\times\Sigma$ for some codimension 1 submanifold $\Sigma$ without boundary, with $\mathbb{R}$ denoting the time direction. In this setting, the coordinate on $\mathbb{R}$ will be denoted with the index 0 and the coordinates on $\Sigma$ with the index $i$, so we write $j_{X}=j^{0}_{X}\mathrm{Vol}_{0}+\sum^{\dim\Sigma}_{i=1}j^{i}_{X}\mathrm{Vol}_{i}$. We continue to focus on the Euclidean and Minkowski signatures and identify $\mathrm{Vol}_{\Sigma}$ with the volume form on $\Sigma$.
@@
Consider $E=M\times\mathbb{R}\to M$, with $M=\mathbb{R}^{n}$ with the Euclidean or Minkowski metric and a first-order Lagrangian $L=L(u,u_{i})$ which is translation invariant. This symmetry can be encoded with $\iota_{X_{\infty}}(\mathrm{d}x^{i})=\epsilon^{i}$, $\iota_{X_{\infty}}(\mathrm{d}u_{I})=0$ so that $\iota_{X_{V}}(\mathrm{d}_{V}u_{I})=-\epsilon^{i}u_{iI}$. Then
@@
$P_{i}[\phi]=\int_{\Sigma}T_{0i}[\phi]\,\mathrm{d}^{n-1}x=\int_{\Sigma}\partial_{0}\phi\partial_{i}\phi\,\mathrm{d}^{n-1}x.$
In particular, the energy for the Euclidean and Minkowski metrics, with $g_{00}=1$, $g_{ii}=\pm 1$, respectively, is
@@
Focusing now on the (1+1)-case for simplicity and choosing a plane wave solution $\phi(t,x)=A\cos(\omega t+g_{11}\kappa x)$ solving the Klein-Gordon equation $(\Delta+m^{2})\phi=0$ with the dispersion relations $m^{2}=\omega^{2}+g_{11}\kappa^{2}$, one recovers that the expected energy and momentum densities (the integrands of $E[\phi]$ and $P_{x}[\phi]$)
@@
and choosing the normalization $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ of the field $\phi$ we then recover the familiar $\langle\mathcal{E}_{dens}[\phi]\rangle^{2}+g_{11}\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}=m^{2}$, indicating that the field $\phi$ has the average density of ‘one particle of mass $m$ per unit volume’, with $\langle\mathcal{E}_{dens}[\phi]\rangle=\omega$ and $\langle\mathcal{P}_{x,dens}[\phi]\rangle=\kappa$. Note that the usual normalization $\sqrt{1/(2\langle\mathcal{E}_{dens}[\phi]\rangle)}$ found when quantizing the field via the CCR is recovered if we write the field $\phi(t,x)=(A/2)(e^{{\mathfrak{i}}(\omega t+g_{11}\kappa x)}+e^{-{\mathfrak{i}}(\omega t+g_{11}\kappa x)})$ in terms of exponentials.
@@
Note that $\mathcal{E}_{dens}[\phi,\phi^{*}]$, $\mathcal{P}_{x,dens}[\phi,\phi^{*}]$ are double the values of the real scalar field theory in Example 2.6 due to the existence of two different modes $\phi$, $\phi^{*}$ in this theory. In this case we find the relation
@@ 3. Variational calculus on the integer lattice $\mathbb{Z}$
We now follow the exact same ideology but replacing the base manifold $M$ with $\mathbb{Z}$ as a discrete noncommutative geometry. This is an application of a general picture whereby $\Omega^{1}$ on a discrete set is precisely a graph, and here we consider $\mathbb{Z}$ as a graph, namely a one-dimensional lattice. Using that $\mathbb{Z}$ is a group, it is convenient to write the calculus in terms of a basis of invariant 1-forms $e^{a}=e^{\pm}$ over the algebra of real valued functions $C(\mathbb{Z})$. We will denote the algebra of functions in the base by $C(\mathbb{Z})$ instead of using $\mathbb{C}(\mathbb{Z})$ since we do not want to necessarily fix the field to be $\mathbb{C}$. In most of the chapter, we will work in fact with real-valued functions but we will also consider complex valued fields in some places, cf. Section 5.3. We do not address the issue of *-structures which would be needed for a full understanding of the noncommutative geometry, since we do not yet have a theory of *-structures on jet bundles, as noted in [21].
The exterior derivative $\mathrm{d}\colon C(\mathbb{Z})\to\Omega^{1}$ is then
@@ 3.1. Construction of the double complex $\Omega(J^{\infty})$
with coordinates $(i,u_{I})$. The jet prolongation map $j_{\infty}\colon C(\mathbb{Z})\to\Gamma(J^{\infty})$ is
@@
$e_{\infty}^{*}(\Phi)(i,\phi)=\Phi(i,(\partial_{I}\phi)(i)),\quad e_{\infty}^{*}(e^{\pm})=\sum_{i}e_{\infty}^{*}(\delta_{i}\mathrm{d}\delta_{i\pm 1})=e^{\pm},$
@@
and recognising the last term, we obtain
@@
where the second relation is redundant, i.e. can be proved by application of the other relations. We used that the $e^{a}$ anticommute to obtain it in this form. Next, using (6) and assuming $\mathrm{d}_{H}$ is a derivation as it should be as inherited from $\Omega(\mathbb{Z}\times F)$ and that $\mathrm{d}_{H}e^{a}=\mathrm{d}e^{a}=0$, we have
@@
$\mathrm{d}_{V}u_{I},\Phi]=0,\quad\{e^{a},\mathrm{d}_{V}u_{I}\}+\mathrm{d}_{V}u_{aI}\wedge e^{a}=0,\quad\{\mathrm{d}_{V}u_{I},\mathrm{d}_{V}u_{J}\}=0,$
@@
$[\mathrm{d}_{V}u_{I},\phi]$ $=[\mathrm{d}u_{I},\phi]-\sum_{a}u_{aI}[e^{a},\phi]=[\mathrm{d}u_{I},\phi]-\sum_{a}u_{aI}(\partial_{a}\phi)e^{a}=0,$
$[\mathrm{d}_{V}u_{I},u_{J}]$ $=[\mathrm{d}u_{I},u_{J}]-\sum_{a}u_{aI}[e^{a},u_{J}]=0,$
@@ 3.2. Euler-Lagrange equations
$R_{a}\colon\Omega^{k}(J^{\infty})\to\Omega^{k}(J^{\infty}),\qquad D_{a}\colon\Omega^{k}(J^{\infty})\to\Omega^{k}(J^{\infty}),$
@@
(11) $R_{a}(\omega\wedge\eta)=R_{a}(\omega)\wedge R_{a}(\eta),\qquad D_{a}(\omega\wedge\eta)=D_{a}(\omega)\wedge R_{a}\eta+\omega\wedge D_{a}(\eta).$
@@
$EL=\left(\frac{\partial L}{\partial u}+\sum_{a}D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\right)\mathrm{d}_{V}u\wedge\mathrm{Vol},\qquad\Theta=\sum_{a}R_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\wedge\mathrm{Vol}_{a}.$
@@
$\mathrm{d}_{V}(L\mathrm{Vol})=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\sum_{a}\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u_{a}\right)\wedge\mathrm{Vol}=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\sum_{a}\frac{\partial L}{\partial u_{a}}D_{a}\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}$
@@
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\sum_{a}D_{a}\left(\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u\right)-\sum_{a}D_{a}\left(\frac{\partial L}{\partial u_{a}}\right)R_{a}\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}$
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\sum_{a}D_{a}\left(\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u\right)+\sum_{a}R_{a}\left(D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)\right)\wedge\mathrm{Vol}$
@@
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\sum_{a}D_{a}\left(\frac{\partial L}{\partial u_{a}}\mathrm{d}_{V}u\right)+\sum_{a}D_{a}\left(D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)+\sum_{a}D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}.$
@@
$=\left(\frac{\partial L}{\partial u}\mathrm{d}_{V}u+\sum_{a}D_{a}\left(R_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)+\sum_{a}D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\right)\wedge\mathrm{Vol}.$
@@
$\mathrm{d}_{V}(L\mathrm{Vol})=\left(\frac{\partial L}{\partial u}+\sum_{a}D_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\right)\mathrm{d}_{V}u\wedge\mathrm{Vol}-\mathrm{d}_{H}\left(\sum_{a}R_{a^{-1}}\left(\frac{\partial L}{\partial u_{a}}\right)\mathrm{d}_{V}u\wedge\mathrm{Vol}_{a}\right).$
@@
As the bundle $\mathbb{Z}\times\mathbb{R}\to\mathbb{Z}$ models a point particle moving in $\mathbb{R}$ evolving in discrete time, we want to reproduce Example 2.1 in this setting, starting with a free particle $q=\phi:\mathbb{Z}\to\mathbb{R}$ with $V(q)=0$. Let $(e^{a},e^{b})=g^{ab}=\delta^{a,b^{-1}}$ be the inverse Euclidean metric on $\mathbb{Z}$, and define the integral on $\mathbb{Z}$ as $\int_{\mathbb{Z}}f\mathrm{Vol}=\sum_{i}f_{i}$. Then in analogy to Example 2.1 we set the action to be
@@
From the general discussion in Section 2.3 we take the key requirement of an interior product along a suitable vector field encoding the translation symmetry and we specify this directly as a map $\iota_{\epsilon}\colon\Omega^{\bullet}(J^{\infty})\to\Omega^{\bullet-1}(J^{\infty})$ by defining it on $\Lambda_{J^{\infty}}$ and extending as a left module map to $\Omega(J^{\infty})=C(J^{\infty})\Lambda_{J^{\infty}}$. Following the treatment for $M$ in Example 2.6, we let
@@
$=\sum_{b}\epsilon^{b}u_{abI}e^{a}=\sum_{b}\epsilon^{b}u_{baI}e^{a}=-\iota_{V}(\mathrm{d}_{V}u_{aI}e^{a}),$
@@
(12) $i_{H}({\rm Vol})=i_{H}(e^{+}\wedge e^{-})=\epsilon^{+}e^{-}-e^{+}\epsilon^{-}=\sum_{a}\epsilon^{a}{\rm Vol}_{a}.$
@@
where (,)(\ ,\ ) is the metric inner product. In our case $\nabla$ obeys $\nabla e^{a}=0$, the Leibniz rule and extends to tensor products in an obvious way (it is a bimodule connection in the sense used in [2] with the generalised braiding given by flip on the basis). Hence divergence-free just amounts to $\sum_{a}\partial_{a^{-1}}(T_{ab})=0$ for all $b$.
@@
$\sum_{a}D_{a^{-1}}T_{ab^{-1}}=-{m\over 4}\sum_{a}\left(2u_{aa^{-1}}+D_{b^{-1}}\left(u_{aa^{-1}}\right)\right)u_{b^{-1}}.$
We therefore see that the last term in the changed expression for $\sum_{a}D_{a^{-1}}T_{ab}$ can be written in terms of $\sum_{a}D_{a^{-1}}T_{ab^{-1}}$, leading to
@@
where we have used $D_{a}=-D_{a^{-1}}R_{a}$, and directly see that the term in brackets is constant along the discrete time axis $\mathbb{Z}$. Expanding it out gives
@@
$E[q]=-{m\over 2}(\partial_{+}q)(\partial_{-}q).$
@@
The current from Proposition 3.4 is conserved when a quadratic potential $V(q)=\mu q^{2}$ is included in the Lagrangian $L(u,u_{a})={m\over 4}\sum_{a}u^{2}_{a}-V(u)$.
@@
$\mathrm{d}_{H}j=-{m\over 4}\sum_{a,b}\epsilon^{b}\left(2u_{aa^{-1}}u_{b}+D_{b}(u_{aa^{-1}})u_{b}\right)+\sum_{b}\epsilon^{b}(D_{b}V(u))\mathrm{Vol}$